# Concept-Aware LoRA for Domain-Aligned Segmentation Dataset Generation

### v1

@@ A Appendix
[…]
B Additional Analysis of Concept Sensitivity
B.1 The Implicit Bias of Gradients Across Layers
Observation (Fig. 6 (a))
We compute the sensitivity scores for each layer by using the norm of the gradient. However, the gradient norm cannot be uniformly scaled across different head types (Q, K, V, OUT), attention types (self, cross), and layers (shallow, deep). To address this, we construct a base gradient to scale the concept loss gradient by referencing the gradient of the original diffusion loss, as described in Section 3.2. We visualize the gradients of both the concept losses and the original diffusion loss in Fig. 6. In this visualization, we separate the self-attention and cross-attention layers to provide clearer distinctions, which differs from the approach in the main paper.
Normalizing Gradients (Fig. 6 (b), Fig. 7)
Therefore, we normalize the gradients using the gradients calculated from the original diffusion loss, as discussed in Section 3.2 and shown in Fig. 3. As shown in Fig. 6 (a), the gradients calculated from the style concept loss and viewpoint concept loss are similar. However, the gradient increase ratio can differ significantly, as illustrated in Fig. 6 (b). The final results of the concept sensitivity are visualized in Fig. 7
Figure 6: The implicit bias of the gradients across the layers (a). Concept sensitivity is calculated by normalizing each gradient with the gradient of the original diffusion loss (b). We highlight style- and viewpoint-sensitive weights in red and blue, respectively. Each column represents the layer position, ranging from shallow to deep, while each row corresponds to the multi-head projection layers for query, key, value, and output in the attention module.
Figure 7: The integrated concept sensitivity of the style- and viewpoint concept sensitivity which we finally used for all experiments.
B.2 Concept Sensitivity According to the Noise Timestep
@@
The amount of added noise, defined by the timestep $t$, is critical hyperparameter for calculating concept sensitivity. We conduct comprehensive experiments to evaluate concept sensitivity in relation to the noise timestep. Visualizations of concept sensitivity across different noise timesteps, along with qualitative and quantitative results, are provided. These experimental findings offer valuable insights into the behavior of concept sensitivity.
Visualization (Fig. 8)
@@
According to our experiment, calculating concept sensitivity at large timesteps (noisy images) does not yield meaningful information about concept sensitivity. For example, style and viewpoint sensitivities appear similar when the timestep is set to 481 out of 1000, as shown in the first column of Fig. 8. This occurs because concept-sensitive layers are less responsive to noisy inputs, which have a high potential to generate any image. Conversely, extremely small timesteps (e.g., 1) also fail to capture concept sensitivity, as the loss from almost clean images does not provide sufficient generative information. Therefore, we explored intermediate timesteps (e.g., 201, 81) and found that the 81st timestep reveals distinct concept sensitivities for style and viewpoint.
Figure 8: Visualizing concept sensitivity across different noise timesteps (481, 201, 81, and 1) shows that the 81st timestep stands out with a significantly distinct concept sensitivity score between style and viewpoint sensitivity compared to the other timesteps.
Qualitative Results (Fig. 9)
@@
Additionally, we fine-tuned 2% of the selected ratio using each concept sensitivity and generated images to qualitatively compare results across different noise timesteps. As shown in Fig. 9, the images generated using intermediate timesteps (201, 81) better align with the intended style and viewpoint.
@@
Finally, we quantitatively compared the intermediate timesteps using the image domain alignment metric, CMMD ($\downarrow$[22], to evaluate the 201st and 81st timesteps. The results indicate that the 81st timestep is the most effective for measuring concept sensitivity, as shown in Table 5. While our approach selects a single timestep to measure concept sensitivity, averaging multiple timesteps could improve the precision and robustness of concept sensitivity, which may be a promising direction for future research.
Extracted Timestep 481 201 81 1
Style 1.920 1.556 1.420 2.383
Viewpoint 1.626 2.132 2.313 2.555
@@
Table 5: The CMMD ($\downarrow$) [22] of the 2% concept-CA-LoRA (style, viewpoint) is evaluated across the extracted timesteps.
Figure 9: According to the various noise timestep, 81st timestep represents the best concept sensitivity, qualitatively. The style of the generated images by Style CA-LoRA is well-aligned, while the generated images by Viewpoint CA-LoRA contain diverse styles.
Figure 10: Measured concept sensitivity according to the various prompt augmentation. The highlighted concept-sensitive layers for each concept (style and viewpoint) remained largely consistent regardless of the prompt augmentation, demonstrating the robustness of concept sensitivity to variations in prompt design.
B.3 Concept Sensitivity According to the Prompt Augmentation
We conduct an additional analysis of the robustness of defining desired concepts based on prompt augmentation. As demonstrated in Section 3.2, we select several hand-crafted prompt augmentations tailored to the desired concept. While this method is flexible and can be generalized to different concepts, it may introduce vulnerabilities in prompt augmentation. To assess its robustness, we measure sensitivity across various prompt augmentations. Specifically, we provide five prompt augmentations for each desired concept (style, viewpoint) derived from the original prompts, as shown below.
[…]
We then calculate the style and viewpoint sensitivity for each prompt augmentation, as shown in Fig. 10. As illustrated in the figure, our proposed method consistently demonstrates high sensitivity to similar regions across all prompt augmentations for styles. Similarly, for viewpoints, augmentations such as top-down, high-angle, and low-angle were applied, and the results indicate that our method highlights similar regions regardless of the specific viewpoint prompt. Based on these findings, we manually select the first three prompts for each desired concept. Nevertheless, developing an automated approach to search for prompt augmentations by leveraging recent LLMs could be a promising direction for enhancing concept sensitivity.
Figure 11: In-domain few-shot semantic segmentation comparison (0.3% Cityscapes) with the hand-crafted layer selection approaches and the variants of CA-LoRA that partition the base weights differently.
Figure 12: Qualitative comparison with the hand-crafted layer selection approaches by generating a “bus” class. While hand-crafted layer selection approaches, and Original LoRA generate similar bus images with the training bus images, we can generate diverse bus images with well-aligned label maps.
C Additional Analysis of Concept-Aware LoRA
In this section, we evaluate the effectiveness of our sensitive weight identification method in capturing the desired concepts by comparing its performance against variants of hand-crafted layer selection methods. Additionally, we show the performance of varients of CA-LoRA by changing the granularity of basic component differently (e.g., block, layer). Specifically, we compare our projection-wise CA-LoRA with block-wise, layer-wise, and integrated projection-wise approaches by analyzing their impact on in-domain semantic segmentation performance.
Experimental setup
The experimental setup follows the in-domain semantic segmentation experiment presented in Section 4.2. To enhance performance, we generate a segmentation dataset using 0.3% of the labeled Cityscapes dataset [9]. For the hand-crafted manual selection baselines, we include LoRA fine-tuning methods that selectively update all self-attention and cross-attention layers, respectively. Since Stable Diffusion XL contains an equal number of self-attention and cross-attention layers, each hand-crafted selection method fine-tunes 50% of the total layers.
To ensure a comprehensive comparison, we also evaluate “DatasetDM”, which utilizes the pretrained model without fine-tuning, and “Original LoRA”, which applies LoRA fine-tuning to all attention layers. Additionally, we explore the granularity of CA-LoRA while maintaining the same concept awareness across block-wise, layer-wise, integrated projection-wise, and projection-wise approaches. Here, “integrated projection-wise” refers to using a single LoRA matrix for all multi-head attention projections. The detailed implementation of projection-wise CA-LoRA is provided in Section G.2, which may help differentiate it from integrated projection-wise CA-LoRA.
Comparison with Hand-crafted Layer Selection Approaches (Figs. 11 and 12)
As shown in Fig. 11, finetuning only the self- or cross-attention layers outperform DatasetDM and Original LoRA. However, their performance does not reach the level of our Style CA-LoRA, which specifically targets the Style-Sensitive Layers in Cityscapes. To analyze this, we provide examples of bus samples generated using the prompt “photorealistic first-person urban street view with bus.” across LoRA finetuning methods for hand-crafted layer selections (self- and cross-attention layers), Original LoRA, and our Style CA-LoRA. As illustrated in Fig. 12, the images of buses seen during training are limited to the top two examples. In the case of hand-crafted layer selection approaches, and Original LoRA, the generated buses closely resemble those seen during training, showing minimal variation. In contrast, our Style CA-LoRA, which selectively fine-tunes only the Style-Sensitive Layers, is capable of generating a diverse range of buses while maintaining the Cityscapes style. We suppose that the diversity in the generated dataset of our method significantly contributed to the superior final performance improvements in semantic segmentation.
Granularity of the Concept-Aware LoRA (Fig. 11)
@@ C Additional Analysis of Concept-Aware LoRA
We explore the performance of CA-LoRA variants that partition the base weights differently. As shown in Fig. 11, these variants generally outperform the baselines. Among them, projection-wise CA-LoRA achieves the highest performance compared to other variants. This suggests that increasing the granularity of the base partition generally enhances performance, as it allows for more fine-grained selection of sensitive weights.
D In-domain Experiments for the General Domain Dataset (Pascal-VOC)
D.1 Experimental setup
In this experiment, we trained on a total of 100 real image-label pairs and evaluated the model using the 1,449 images in the Pascal-VOC validation set. For the text-to-image generation model, we applied the same style sensitivity score used in the Cityscapes experiment, setting the selected proportion to 10%. During the training of the text-to-image generation model, the prompt ”a photo” was used. For training the label generator and generating the dataset, the prompt ”a photo of a {class names}” was employed. The label generator was trained with a batch size of 4 for 90K iterations, ultimately producing 2,000 image-label pairs. When utilizing the generated dataset, the process was consistent with the in-domain semantic segmentation experiments. Specifically, Mask2Former was trained on the real dataset for 90K iterations (Baseline), followed by fine-tuning on the combined real and generated dataset for an additional 30K iterations. Additionally, we include an additional fine-tuned baseline (Baseline (FT)) that is solely fine-tuned on the same real dataset for a fair comparison in terms of the total iterations. All other hyper-parameters remained identical to those used in the Cityscapes in-domain semantic segmentation experiment, as detailed in Tables 13, 14 and 15.
D.2 Quantitative and Qualitative Results
As shown in Table 6, using Style CA-LoRA on the Pascal-VOC dataset resulted in a performance improvement of 0.93 mIoU. In contrast, DatasetDM, which omitted the fine-tuning process for the text-to-image generation model, showed a performance drop of 8.43 mIoU. This highlights the importance of concept-aware finetuning for style, even in general datasets beyond urban-scene datasets. Fig. 13 provides further insight into the role of style information. A significant image domain gap is evident between the images generated by the pretrained text-to-image generation model and the dataset generated using Pascal-VOC. This demonstrates the impact of image domain alignment. Quantitatively, the CMMD, which was 1.46 for the pretrained model, decreased to 0.81 after alignment, illustrating the reduced domain gap and its contribution to performance improvement.
Method Training Dataset Total Iterations Segmentation Performance (mIoU)
# Real # Generated
Baseline 100 90K 44.59
For a fair comparison, we fine-tune the baseline
for an additional 30K iterations using real or generated datasets.
Baseline (FT) 100 120K 44.39 (-0.20)
DatasetDM 100 2,000 120K 36.16 (-8.43)
Ours 100 2,000 120K 45.52 (+0.93)
Table 6: In-domain segmentation performance (mIoU) of the Pascal-VOC dataset in the few-shot setting (100 image-label pairs). In the first row, we trained Mask2Former on various fractions of the Cityscapes dataset (Baseline). Then, we fine-tuned the baseline on DatasetDM and our generated datasets with 30K iterations and evaluated the performance of the fine-tuned segmentation models. Additionally, we include an additional fine-tuned baseline (Baseline (FT)) that is solely fine-tuned on the same real dataset for a fair comparison in terms of the total iterations.
E Additional Qualitative and Quantitative Results
E.1 Comparison of Image-Label Alignment
Quantitative Comparison (Table 7)
Since the generated images lack ground-truth label maps, we measure image-label alignment using predictions from a pretrained segmentor. Specifically, we use the predictions from the pretrained Mask2Former model [^5] https://github.com/facebookresearch/Mask2Former, which was fully supervised on the 100% Cityscapes dataset and achieves a 79.40 mIoU, as a proxy for the ground truth mask. Since this method is valid only when the pre-trained segmentor significantly outperforms the label generator, we conduct the image-label alignment experiment in a 0.3% few-shot setting.
Analysis of the Qualitative Comparison (Fig. 5)
We compare not only image quality but also image-label alignment across DatasetDM [51], original LoRA, and our Viewpoint- and Style CA-LoRA in the 0.3% few-shot segmentation setting. As shown in the generated labels in Fig. 5, DatasetDM fails to generate corresponding labels, while our Style CA-LoRA generates high-quality corresponding labels. We suppose the reason is grounded by the domain gap between the pretrained T2I model (SDXL [37]) and the source dataset (Cityscapes [9]). As mentioned Section 3.4, DatasetDM trains a label generator using the images of the source domain (e.g., Cityscape images) without performing domain adaptation of the pre-trained T2I model. In other words, the domain gap exists between the label generator and the text-to-image model since the label generator is updated with the Cityscapes images while the original pretrained text-to-image model is not. Due to this domain gap, the intermediate features extracted from the original text-to-image model often fail to reflect the knowledge required for generating label maps of Cityscapes when used as input for the label generator. On the other hand, Style CA-LoRA effectively adapts the T2I model to generate Cityscapes-style images. Therefore, Style CA-LoRA can generate high-quality labels by reducing the domain gap between the intermediate features. However, although the image-label alignment has increased according to the increasing proportions of the selected layers, it does not always provide a better dataset, as shown in our ablation study Table 12 due to the image memorization problem.
Desired Concept Proportion of T2I Model Layers for Concept-Aware Fine-tuning
0% (Pretrained) 1% 2% 3% 5% 10% 100% (Original LoRA)
Style 25.18 37.01 39.37 40.94 47.84 48.86 60.10
Viewpoint 33.27 29.42 31.44 34.94 38.28
Table 7: Image-Label Alignment (mIoU) ($\uparrow$) across the segmentation dataset generation approaches in a few-shot setting (0.3%).
E.2 Additional Analysis of Our Generated Dataset on the Domain Generalization Setting
In this section, we aim to compare and analyze the performance of the Viewpoint CA-LoRA against other baselines that have been applied to urban-scene segmentation in domain generalization. This analysis comprises qualitative assessments (Fig. 14) alongside quantitative evaluations of image domain alignment (Table 8) and image-label alignment (Table 10), similar to Section 4.3 and Section E.1, respectively.
Image Domain Alignment (Tables 8 and 14)
For domain generalization in urban-scene segmentation, we generated urban-scene images under various adverse weather conditions (e.g., “foggy”, “night-time”, “rainy”, and “snowy”). In this section, we assess the domain gap between our generated adverse weather conditions and the real ACDC [45] dataset. Qualitatively, as shown in Fig. 14, our approach generates images that are more realistic and better aligned compared to DatasetDM and InstructPix2Pix, which rely on pretrained models without fine-tuning. When compared to DATUM [4], our method achieves a similar level of image domain alignment while generating more diverse scenes. Quantitatively, we used CMMD [22] to measure image domain alignment, and the results are presented in Table 8. These results show that the proposed generated dataset demonstrates a significant performance gap over DatasetDM and InstructPix2Pix. More importantly, it achieves competitive performance with DATUM, which requires training individual models separately for each weather condition using a target domain image from the ACDC dataset.
@@ E.2 Additional Analysis of Our Generated Dataset on the Domain Generalization Setting
Table 8: Comparison of image domain alignment with image generation baselines on four adverse weather conditions. The alignment is measured between the generated images and the ACDC dataset for each weather condition (CMMD $\downarrow$). †DATUM trained 4 models for each weather condition in the ACDC dataset, using an additionally provided single target domain image per condition.
Figure 14: Qualitative results for generating image-label pairs in domain generalization settings. The proposed approach demonstrates its efficacy in both image domain alignment and image-label alignment.
@@
Table 9: Segmentation performance of the Pretrained and Finetuned Mask2Former (M2F) [6] on the adverse weather condition dataset (ACDC [45]). Starting with the pretrained M2F model trained on the Cityscapes dataset [9], we further fine-tuned the model on the ACDC training set for each individual weather condition (learning rate is 3e-6, the batch size is 2, and the number of iterations is 30K). This approach resulted in highly effective segmentation models tailored to specific weather conditions, serving as pseudo ground-truth masks for evaluating image-label alignment in domain generalization settings.
Method Foggy Night-time Rainy Snowy Average
@@
Table 10: Comparison of image-label alignment with baselines. While InstructPix2Pix provides reliable image-label alignment by fixing Cityscapes labels and applying style transfer only to the weather conditions of the images, its ability to generate diverse scenes is constrained by the fixed labels. In contrast, when comparing methods that generate labels, our approach demonstrates better image-label alignment than DatasetDM.
Figure 15: (a) Class-wise performance improvements (IoU) and (b) Label proportions for the original Cityscapes, Ours, and “Ours (+ More Person Samples)”. “Ours (+ More Person Samples)” includes an additional 500 samples for the “person” class to balance the label proportions. As shown in the class-wise performance, significant improvements were achieved for rare classes such as “bus”, “fence”, and “bicycle”, as highlighted by the blue dotted lines. While some classes, such as “person” and “rider”, showed degradation (indicated by red dotted lines), this was due to the lower number of generated samples for these classes. By generating additional samples for these specific classes, a more balanced performance improvement can be achieved, ultimately increasing the overall average performance.
Image-Label Alignment (Figs. 14, 9 and 10)
We compare image-label alignment to evaluate how reliably label maps are generated for datasets aimed at domain generalization. Qualitatively, as shown in Fig. 14, DATUM generates only images and sets them as pseudo-target domains to apply UDA methods, meaning that no labels are generated. In the case of InstructPix2Pix, style transfer is performed on Cityscapes image-label pairs, using the labels from Cityscapes directly. While this ensures high-quality labels, severe editing can occasionally cause alignment issues, as seen in the foggy examples. Finally, comparing DatasetDM and our method, both of which generate labels directly, shows that our approach achieves significantly better label generation quality compared to DatasetDM.
We then proceed to evaluate image-label alignment quantitatively. As discussed in Section E.1, the generated images lack actual ground truth for domain generalization datasets. Therefore, we rely on pseudo ground truth generated by a highly accurate segmentor. Since no off-the-shelf urban-scene semantic segmentation model consistently performs well across diverse domains, we took several steps to develop a more reliable segmentor. First, as described in Section E.1, we began with a pretrained Mask2Former (M2F) model trained on the full Cityscapes dataset. However, as shown in Table 9, this model is susceptible to adverse weather conditions. To address this limitation, we fine-tuned the pretrained M2F model individually for each of the four adverse weather conditions in the ACDC training set. Since this dataset is not accessible to DatasetDM or our method, the fine-tuned M2F models are guaranteed to outperform those methods. The specific performance improvements on the ACDC validation set are detailed in Table 9.
The results of measuring image-label alignment using the fine-tuned M2F models are presented in Table 10. As shown in the table, InstructPix2Pix, which directly uses Cityscapes labels and only slightly edits the weather conditions of the images, demonstrates an advantage in image-label alignment. Despite its high image-label alignment, we highlighted the limited performance improvements of InstructPix2Pix in Table 2 and Section 4.2, attributing this to the lack of scene diversity caused by its reliance on fixed segmentation label maps. When comparing methods that generate labels, our approach achieves better image-label alignment than DatasetDM. This improvement stems from our text-to-image generation model learning viewpoints from Cityscapes. As a result, even with the same label generator architecture, our finetuned text-to-image generation model provides representations with a smaller domain gap when generating images based on the Cityscapes dataset.
E.3 Additional Qualitative Results
Additional examples (Figs. 20, LABEL:, 21, LABEL: and 22)
We illustrate the changes in generated images using Style- and Viewpoint CA-LoRA as the proportion of selected layers varies (1%, 2%, 3%, 5%, and 10%). As shown in Fig. 20, both CA-LoRAs effectively focus on the target concept with a small proportion of selected layers. However, as the proportion increases, other concepts are gradually learned, as demonstrated in the 10% layer selection. For example, the Viewpoint CA-LoRA shows a slight adaptation to the Cityscapes style. This flexibility allows for manual adjustment of the extent to which other concepts are learned, depending on the specific problem settings or datasets.
Furthermore, we provide additional examples of the generated datasets used to improve segmentation performance in fully supervised and domain generalization settings. The additional examples of our generated datasets are available in Figs. 21 and 22.

### v2

@@ B.2 Concept Awareness per Prompt Augmentation
We then calculate the style and viewpoint awareness for each prompt augmentation, as shown in Figure 10. As illustrated in the figure, our proposed method consistently demonstrates high awareness to similar regions across all prompt augmentations for styles. Similarly, for viewpoints, augmentations such as top-down, high-angle, and low-angle were applied, and the results indicate that our method highlights similar regions regardless of the specific viewpoint prompt. Based on these findings, we manually select the first three prompts for each desired concept. Nevertheless, developing an automated approach to search for prompt augmentations by leveraging recent LLMs could be a promising direction for enhancing concept awareness.
B.3 Concept Awareness per Noise Timestep
@@
The amount of added noise, defined by the timestep $t$, is critical hyperparameter for calculating concept awareness. We conduct comprehensive experiments to evaluate concept awareness in relation to the noise timestep. Visualizations of concept awareness across different noise timesteps, along with qualitative and quantitative results, are provided. These experimental findings offer valuable insights into the behavior of concept awareness.
Visualization (Figure 11)
@@
According to our experiment, calculating concept awareness at large timesteps (noisy images) does not yield meaningful information about concept awareness. For example, style and viewpoint sensitivities appear similar when the timestep is set to 481 out of 1000, as shown in the first column of Figure 11. This occurs because concept-aware layers are less responsive to noisy inputs, which have a high potential to generate any image. Conversely, extremely small timesteps (e.g., 1) also fail to capture concept awareness, as the loss from almost clean images does not provide sufficient generative information. Therefore, we explored intermediate timesteps (e.g., 201, 81) and found that the 81st timestep reveals distinct concept sensitivities for style and viewpoint.
Qualitative Results (Figure 12)
@@
Additionally, we fine-tuned 2% of the selected ratio using each concept awareness and generated images to qualitatively compare results across different noise timesteps. As shown in Figure 12, the images generated using intermediate timesteps (201, 81) better align with the intended style and viewpoint.
@@
Finally, we quantitatively compared the intermediate timesteps using the image domain alignment metric, CMMD ($\downarrow$[21], to evaluate the 201st and 81st timesteps. The results indicate that the 81st timestep is the most effective for measuring concept awareness, as shown in Table 5. While our approach selects a single timestep to measure concept awareness, averaging multiple timesteps could improve the precision and robustness of concept awareness, which may be a promising direction for future research.
@@
Table 5: Timestep ablation. The CMMD ($\downarrow$) and final segmentation performance of the CA-LoRA across the extracted timesteps.
Timestep CMMD ($\downarrow$) Performance ($\uparrow$)
1 2.383 40.49
81 1.420 44.13
201 1.556 41.68
481 1.920 40.80
B.4 Layer selection strategy of CA-LoRA
Grouping Granularity of CA-LoRA (Figs. 13 and 6)
@@ B.4 Layer selection strategy of CA-LoRA
As illustrated in Figure 13, we explore the performance of CA-LoRA variants that partition the base weights differently. As shown in Tab. 6, these variants generally outperform the baselines. Among them, projection-wise CA-LoRA achieves the highest performance compared to other variants. This suggests that increasing the granularity of the base partition generally enhances performance, as it allows for more fine-grained selection of sensitive weights.
Figure 13: Visualizing the concept awareness (style) according to various grouping granularity.
Table 6: Granularity ablation. Number of trainable parameters and final segmentation performance by changing grouping granularity. Only 2% of projection-wise parameters can achieve the best results across all grouping granularity.
Grouping Granularity % Params. Performance ($\uparrow$)
No Finetuning (DatasetDM) 0% 42.82
Block-wise 32% 43.67 (+0.85)
Layer-wise 4.8% 43.26 (+0.44)
Integrated Projection-wise 5% 43.63 (+0.81)
Projection-wise (Ours) 2% 44.13 (+1.31)
Figure 14: Qualitative results of Style- and Viewpoint CA-LoRA across different layer proportions (1%, 2%, 3%, 5%, and 10%). Selecting only 2% of the layers is already sufficient to learn the target concepts, while larger layer proportions increasingly introduce unintended concept leakage.
Layer Proportions of CA-LoRA (Figure 14)
We analyze the qualitative behavior of Style CA-LoRA and Viewpoint CA-LoRA when varying the proportion of selected layers (1%, 2%, 3%, 5%, and 10%). The results show that selecting only 2% of the layers is sufficient to learn the intended concepts: Style CA-LoRA produces images that accurately reflect the target style, whereas Viewpoint CA-LoRA generates images aligned with the desired viewpoint. However, as the layer proportion increases, the model gradually begins to encode unintended, non-target concepts, indicating growing interference between concepts at higher proportions.
Appendix C CA-LoRA for Domain Generalization
In this section, we comprehensively compare the performance of CA-LoRA with baseline methods used for domain-generalized urban-scene segmentation. We apply both image-domain alignment and image-label alignment when constructing the domain-generalization dataset, following the procedure described in Section 4.3, and also provide the comparison with DGInStyle [22].
Table 7: Image-domain alignment in domain generalization.  DATUM utilizes one provided target-domain image per condition, whereas the other methods do not.
@@ Appendix C CA-LoRA for Domain Generalization
Table 8: Image-label alignment in domain generalization. The first two rows report the segmentation performance of the pretrained and finetuned Mask2Former (M2F) [5] on the ACDC adverse-weather dataset [45]. The following four rows present a comparison of image-label alignment across baseline methods, evaluated using the finetuned M2F model.
@@
DATUM - - - - -
@@
Image Domain Alignment (Table 7)
For domain generalization in urban-scene segmentation, we generated urban-scene images under various adverse weather conditions (e.g., “foggy”, “night-time”, “rainy”, and “snowy”). In this section, we assess the domain gap between our generated adverse weather conditions and the real ACDC [45] dataset. Quantitatively, we used CMMD [21] to measure image domain alignment, and the results are presented in Table 7. These results show that the proposed generated dataset demonstrates a significant performance gap over DatasetDM and InstructPix2Pix. More importantly, it achieves competitive performance with DATUM, which requires training individual models separately for each weather condition using a target domain image from the ACDC dataset.
Image-Label Alignment (Table 8)
We compare image-label alignment to assess how reliably label maps can be produced for domain-generalization datasets. Because the generated images do not have real ground-truth annotations, we instead use pseudo ground-truth produced by a highly accurate segmentation model. As no off-the-shelf urban-scene segmentation model achieves consistently strong performance across diverse domains, we manually finetune a reliable pretrained segmentor (Mask2Former [5]) on the ACDC dataset [45]. The first two rows in the figure show the segmentation performance on the ACDC validation set before and after finetuning, respectively.
The results of measuring image-label alignment using the fine-tuned M2F models are provided in Table 8. As shown in the table, our approach achieves superior image-label alignment compared to DatasetDM also on domain generalization setting. InstructPix2Pix, which directly reuses Cityscapes labels and applies only minor weather edits in image, exhibits an advantage in image-label alignment. However, despite this strong alignment, we previously highlighted its limited performance gains in Tables 1 and 2, which we attribute to the lack of scene diversity inherent to its use of fixed segmentation label maps (see Figure 5).
Table 9: Comparison with DGInStyle of generated datasets for domain generalization of urban-scene segmentation (mIoU). We highlight the performance improvements are larger than DGInStyle, especially on HRDA backbone. *The gray color indicates the reported score from the DGInStyle authors.
DG Method Method ACDC DZ BDD MV Average
ColorAug Baseline 52.38 23.00 53.33 60.06 47.19
DGInStyle 55.19 26.83 55.18 59.95 49.29 (+2.10)
Baseline 53.12 25.69 53.00 59.81 47.91
CA-LoRA 56.07 29.75 54.35 61.40 50.39 (+2.49)
DAFormer Baseline 55.15 28.28 54.19 61.67 49.82
DGInStyle 57.74 28.55 56.26 62.67 51.31 (+1.48)
Baseline 53.98 27.82 54.29 62.69 49.70
CA-LoRA 55.83 31.68 54.68 63.09 51.32 (+1.63)
HRDA Baseline 59.70 31.07 58.49 68.32 54.40
DGInStyle 61.00 32.60 58.84 67.99 55.11 (+0.71)
Baseline 58.48 29.46 56.12 64.27 52.08
CA-LoRA 58.93 34.41 56.56 64.54 53.61 (+1.53)
Comparison with DGInStyle (Table 9)
As shown in Table 9, our CA-LoRA achieves performance gains comparable to, and in most cases larger than, those reported for DGInStyle [22]. As their code and Cityscapes-generated data are unavailable, and their baseline accuracy is marginally higher than ours, we compare only the performance gains rather than the absolute numbers. We will update the comparison once the official code becomes available. Notably, our approach consistently exhibits substantial improvements when applied to the HRDA method. In contrast, DGInStyle generates images using fixed label maps, which limits its ability to construct semantically diverse scenes. We hypothesize that this restricted diversity contributes to the relatively small performance gains observed when DGInStyle is used with advanced DG methods such as HRDA.
Appendix D Experiments in General Domain
D.1 Experimental setup
In this experiment, we trained on a total of 100 real image-label pairs and evaluated the model using the 1,449 images in the Pascal-VOC validation set. For the text-to-image generation model, we applied the same style awareness score used in the Cityscapes experiment, setting the selected proportion to 10%. During the training of the text-to-image generation model, the prompt ”a photo” was used. For training the label generator and generating the dataset, the prompt ”a photo of a {class names}” was employed. The label generator was trained with a batch size of 4 for 90K iterations, ultimately producing 2,000 image-label pairs. When utilizing the generated dataset, the process was consistent with the in-domain semantic segmentation experiments. Specifically, Mask2Former was trained on the real dataset for 90K iterations (Baseline), followed by fine-tuning on the combined real and generated dataset for an additional 30K iterations. All other hyper-parameters remained identical to those used in the Cityscapes in-domain semantic segmentation experiment, as detailed in Tables 12, 13 and 14. Notably, the diffusion timestep ($t=81$) and the real-to-synthetic data ratio ($1:1$) are directly transferred from the Cityscapes experiment without re-tuning, demonstrating the robustness of our key hyperparameters across datasets.
D.2 Quantitative and Qualitative Results
As shown in Table 10, using Style CA-LoRA on the Pascal-VOC dataset resulted in a performance improvement of 0.93 mIoU. In contrast, DatasetDM, which omitted the fine-tuning process for the text-to-image generation model, showed a performance drop of 8.43 mIoU. This highlights the importance of concept-aware finetuning for style, even in general datasets beyond urban-scene datasets. Figure 15 provides further insight into the role of style information. A significant image domain gap is evident between the images generated by the pretrained text-to-image generation model and the dataset generated using Pascal-VOC. This demonstrates the impact of image domain alignment. Quantitatively, the CMMD, which was 1.46 for the pretrained model, decreased to 0.81 after alignment, illustrating the reduced domain gap and its contribution to performance improvement.
Figure 16: Class-wise performance improvements and label proportions. (a) Class-wise IoU improvements and (b) label proportions for Cityscapes, Ours, and Ours (+ More Person Samples). Adding 500 additional “person” samples balances label proportions. Rare classes such as “bus”, “fence”, and “bicycle” show clear gains, while decreases in “person” and “rider” arise from insufficient samples. Supplementing these classes further leads to more balanced improvements and a higher overall average.
Table 10: In-domain segmentation performance (mIoU) of the Pascal VOC dataset. In the first row, we report the performance of Mask2Former trained on different fractions of the PASCAL VOC dataset (Baseline). While DatasetDM often degrades segmentation performance due to incorrect labels, CA-LoRA consistently improves the results.
Method Fraction of the PASCAL VOC Dataset
100 images ($\sim$7%) 1,464 images (100%)
Baseline 44.39 72.54
DatasetDM 36.16 (-8.43) 70.78 (-1.76)
CA-LoRA (Ours) 45.52 (+0.93) 73.72 (+1.18)
Appendix E Class-Specific Dataset Generation